OpenStudy (anonymous):

A snail moving across the lawn for her evening constitutional crawl is attracted to a live wire. on reaching the wire her speed increases at a constant rate&it doubles from 0.001 ms^-1 in 10sec. she remains at this speed for a further 15sec. while she remains contact with the wire. how to calculate the distance traveled during this time??? thanks :)

OpenStudy (anonymous):

Double speed = .002ms^-1 Average rate of change = .002+.001=.003 divided by 2 = .0015. .0015 x 10 = .015 .002 x 5 remaining seconds = .01 .01+.015 = .0225 ms^-1

OpenStudy (anonymous):

constant acceleration equation Distance traveled=(Initial Velocity)*(time) + (acceleration)*((time)^2) find the average acceleration and use that formula over the first 10 secs for the next 15 seconds u can use distance=(velocity)*(time)

OpenStudy (anonymous):

ive never seen this formula before, I DONT KNOW HOW TO USE IT *panic*

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

the distance travelled one :S im not even sure whether ive got the number that id need for it :(

OpenStudy (anonymous):

But the answers are using two different formula.

OpenStudy (anonymous):

Maybe take a look here: http://en.wikipedia.org/wiki/Acceleration

OpenStudy (anonymous):

r u in calculus?

OpenStudy (anonymous):

i dunno, it mechanics 1 motion :S

OpenStudy (anonymous):

have u done integration b4?

OpenStudy (anonymous):

thats dy/dx stuff right ??

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok basically to use the equation i gave u need an acceleration n it needs to b a constant. u multiply this acceleration by the time u r accelerating squared so \[t^2\]

OpenStudy (anonymous):

then u add this to ur initial velocity(the velocity u started out at) multiplied by time so the whole equation is \[d=vt+at^2\]

OpenStudy (anonymous):

v is the velocity u started out at in this case its .001ms^-1 a is the acceleration u have to find. since u know the final acceleration which is .002ms^-1 since it doubled, u can find the average acceleration. (final velocity - initial velocity)/time so: (.002-.001)/10= .001/10= .0001ms^-2 t is the time u travaled in this case its 10 seconds plug all of the numbers in and u will find the distance for the 1st part

OpenStudy (anonymous):

Isn't Jakesmilyface formula equivalent to yours?

OpenStudy (anonymous):

u know, the (u+v)/2

OpenStudy (anonymous):

yea, but he explained and he didnt :)

OpenStudy (anonymous):

for the 2nd part u can use the same exact equation but the acceleration this time is 0 because it stayed at the same speed thus no acceleration so this time the velocity is .002ms^-1 the acceleration is 0 and the time is now 15 seconds plug in all the numbers n u can get the answer. Then add it to ur first answer n u will get the answer u need

OpenStudy (anonymous):

yes it is i think

OpenStudy (anonymous):

I agree, it's a great explanation. I'm just asking if he agrees that the 2 are equivalent.

OpenStudy (anonymous):

??? second answer :S im confused

OpenStudy (anonymous):

but i did what you said and came up with 0.003 but then i didnt use the thing that was just being worked out the 0.0001ms^-2

OpenStudy (anonymous):

its a 2 part problem because at 1point the snail accelerated which means u have to find the distance he traveled while he was accelerating and then u have to find the distance he traveled at his top speed. Think of it as when u r driving a car on the highway when ur accelerating to the speed limit(hopefully no higher) u r still traveling some distance right? u need an equation to solve the amount of distance u traveled then n when u get to ur top speed u need to use a different equation because u r no longer accelerating

OpenStudy (anonymous):

but doesnt that mean that i should have used that 0.0001 to work out the answer?? cause i didnt and got 0.03, is that right??

OpenStudy (anonymous):

I still like to know if you agree that the formulas for part 1 are equivalent?

OpenStudy (anonymous):

the answer should b .05m

OpenStudy (anonymous):

oh :( WHY?!?! why?!?! why cant I do it :s

OpenStudy (anonymous):

ok im gonna work it out n show my work

OpenStudy (anonymous):

thanks you :)

OpenStudy (anonymous):

1st part d=vt+at^2 v=.001 a=(.002-.001)/10=(.001)/10=.0001 t=10 so it when i replace everythin it becomes d=(.001)(10)+(.0001)(10^2)= .01+.01=.02 2nd part d=vt+at^2 v=.002 t=15 a=0 so it becomes d=(.002)(15)= .03 add the 1st n 2nd part together .02+.03=.05 so in that 25 seconds it traveled .05m

OpenStudy (anonymous):

also i looked at jakesmileyface's work n it isnt correct

OpenStudy (anonymous):

accelerating from one speed to another is not the same as driving at the average of the speeds

OpenStudy (anonymous):

thanks your working out really helped me :)

OpenStudy (anonymous):

anytime